OpenStudy (anonymous):

Please help me with these I have no clue what to do. Please I am begging you. 1.) Which of the following is an actual zero of f(x) = x3 + 5x2 - 4x - 20? A) -20 B) -5 C) 1 D) 4 2.) If -nx2 + tx + c = 0, what is x equal to? A) x is equal to minus t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n B) x is equal to t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n C) x is equal to t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n D) x is equal to minus t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n 3.) The graphs of f(x) and g(x) are shown below: graph of function f of x equals x squared minus x minus 12. Graph of function g of x equals negative 2.5 times x minus 2 What are the solutions to the equation f(x) = g(x)? A) x = -3, 4 B) x = -4, 2.5 C) x = -0.8, 2.5 D) x = 8, -8 4.) Which statement best describes the graph of -x3 - x2 + 4x + 4? A) It starts down on the left and goes up on the right and intersects the axis at x = -2, 2, and 4. B) It starts down on the left and goes up on the right and intersects the axis at x = -2, -1, and 2. C) It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. D) It starts up on the left and goes down on the right and intersects the axis at x = -2, 2, and 4. 6.) Which graph shows the quadratic function y = 3x2 + 12x + 14? The following graph is labeled A: A four quadrant graph with a parabola opening up, passing through the points (-3, 1), (-2, -2,) and (-1, 1) with the vertex at(2, -2). The following graph is labeled B: A four quadrant graph with a parabola opening up, passing through the points (1, 4), (2, 1,) and (3, 4) with the vertex at(2, 1). The following graph is labeled C: A four quadrant graph with a parabola opening up, passing through the points (-3, 5), (-2, 2,) and (-1, 5) with the vertex at(-2, 2). The following graph is labeled D: A four quadrant graph with a parabola opening up, passing through the points (1, 1), (2, -2,) and (3, 1) with the vertex at(2, -2). A) A B) B C) C D) D 9.) A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 34t + 80 The average rate of change of f(t) from t = 5 seconds to t = 7 seconds is _____ feet per second. 10.) What are the exact solutions of x2 + 5x + 1 = 0? A) x = the quantity of 5 plus or minus the square root of 21 all over 2 B) x = the quantity of negative 5 plus or minus the square root of 21 all over 2 C) x = the quantity of 5 plus or minus the square root of 29 all over 2 D) x = the quantity of negative 5 plus or minus the square root of 29 all over 2 12.) A company launched four new products. The market price, in dollars, of the four products after different number of years is shown below: Table displaying data. The first column one shows the heading number of years, and rows product 1, product 2, product 3, product 4. The next column heading says 1 with rows 2.00, 2.20, 2.20, and 2.80. The next column header says 2 with rows 4.00, 4.80, 6.40 and 3.20. The next column header says 3 with rows 8.00, 6.80, 11.60, 3.20. The last colunn heading says 4 with rows 16.00, 9.20, 18.80 and 2.80. Based on the data given, the price of which product will eventually exceed all others? A) Product 1 B) Product 2 C) Product 3 D) Product 4 13.)The following function represents the production cost f(x), in dollars, for x number of units produced by company 1: f(x) = 0.05x2 - 7x + 300 The following table represents the production cost g(x), in dollars, for x number of units produced by company 2: x g(x) 0.6 899.58 0.8 899.52 1 899.50 1.2 899.52 1.4 899.58 Based on the given information, the minimum production cost for company _____ is greater. [Put 1 or 2 in the blank space] 14.)A basketball is thrown upwards. The height f(t), in feet, of the basketball at time t, in seconds, is given by the following function: f(t) = -16t2 + 94t + 12 Which of the following is a reasonable domain of the graph of the function when the basketball falls from its maximum height to the ground? A) 3 < t < 6 B) 2 < t < 5 C) 1 < t < 4 D) 0 < t < 3 16.)A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? A) Add 9 to both sides of the equation B) Add 5 to both sides of the equation C) Subtract 5 from both sides of the equation D) Subtract 9 from both sides of the equation 17.)If (x + 1)2 = 55, what is the approximate positive value of x? A) 2.71 B) 5.42 C) 6.42 D) 7.35 19.) The following function shows the relationship between the selling price (s), and profit P(s), in dollars, for a company: P(s) = -20s2 + 1,400s - 12,000 Which statement best describes the intervals where the company's profit increases, decreases, or records a maximum? A) It is least when the selling price is $30. B) It is greatest when the selling price is $30. C) It decreases when the selling price increases beyond $35. D) It increases when the selling price increases beyond $35. 20.)The graphs of f(x) and g(x) are shown below: graph of function f of x open upward and has its vertex at negative 7, 0. Graph of function g of x opens upward and has its vertex at negative 7, negative 2. If f(x) = (x + 7)2, which of the following is g(x) based on the translation? A) g(x) = x2 + 18x + 81 B) g(x) = x2 + 10x + 25 C) g(x) = x2 + 14x + 47 D) g(x) = x2 + 14x + 51

OpenStudy (anonymous):

\[ f(x) = x^3 + 5x^2 - 4x - 20\] is \[f(20)=0\]?

OpenStudy (anonymous):

actually it would be \(f(-20)\) but that is not zero either, so not that one hmmm none of those answers are correct though maybe the first one should have been \(-2\)

OpenStudy (anonymous):

\[f(x) = x^3 + 5x^2 - 4x - 20\] \[f(-2) = (-2)^3 + 5(-2)^2 - 4(-2) - 20\]

OpenStudy (anonymous):

compute that one i bet you get zero

OpenStudy (anonymous):

Ok. Yes I got 0. :) What do I do now?

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

It is not one of the multiple choice answers though.

OpenStudy (anonymous):

1.) Which of the following is an actual zero of f(x) = x3 + 5x2 - 4x - 20? A) -20 B) -5 C) 1 D) 4 then there is a typo it should be A) -2

OpenStudy (anonymous):

none of the others are zeros

OpenStudy (anonymous):

ohhh nvm it is \(-5\)

OpenStudy (anonymous):

Thank you so much! You are a life saver. Can you help me with a couple more please?

OpenStudy (anonymous):

compute \[f(-5)\] and you get zero as well sorry

OpenStudy (anonymous):

you got another question?

OpenStudy (anonymous):

Haha it's ok. And I see what you mean. Thank you so much!! :) :) :)

OpenStudy (anonymous):

2.) If -nx2 + tx + c = 0, what is x equal to? A) x is equal to minus t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n B) x is equal to t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n C) x is equal to t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n D) x is equal to minus t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n

OpenStudy (anonymous):

\[nx^2 + tx + c = 0\] put \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] only in place of \(a\) put \(n\) and in place of \(b\) put \(t\) it will look just the same only with different variables

OpenStudy (anonymous):

i cant read all that muck you wrote above, but it should look like \[x=\frac{-t\pm\sqrt{t^2-4nc}}{2n}\]

OpenStudy (anonymous):

That's exactly what I got! :)

OpenStudy (anonymous):

k good

OpenStudy (anonymous):

Except the multiple choice only has it that it is a -2..:(

OpenStudy (anonymous):

However the closest one would be A. Should I put that?

OpenStudy (anonymous):

hold on i think it looks just like D

OpenStudy (anonymous):

letme check

OpenStudy (anonymous):

oh i made a typo didn't i? it is \(-n\) not \(n\)

OpenStudy (anonymous):

\[-nx^2 + tx + c = 0\]

OpenStudy (anonymous):

ohhhh haha ok. So than it would end up being D?

OpenStudy (anonymous):

\[x=\frac{-t\pm\sqrt{t^2+4an}}{-2n}\]

OpenStudy (anonymous):

you got that choice?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

however there is no an

OpenStudy (anonymous):

if so use it i think it is D but it is too hard for me to read all the words

OpenStudy (anonymous):

right no \(an\) it should be \(cn\)

OpenStudy (anonymous):

\[x=\frac{-t\pm\sqrt{t^2+4cn}}{-2n}\]

OpenStudy (anonymous):

Ok thank you! I have a couple more if that's ok?

OpenStudy (anonymous):

3.) The graphs of f(x) and g(x) are shown below: graph of function f of x equals x squared minus x minus 12. Graph of function g of x equals negative 2.5 times x minus 2 What are the solutions to the equation f(x) = g(x)? A) x = -3, 4 B) x = -4, 2.5 C) x = -0.8, 2.5 D) x = 8, -8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

4.) Which statement best describes the graph of -x3 - x2 + 4x + 4? A) It starts down on the left and goes up on the right and intersects the axis at x = -2, 2, and 4. B) It starts down on the left and goes up on the right and intersects the axis at x = -2, -1, and 2. C) It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. D) It starts up on the left and goes down on the right and intersects the axis at x = -2, 2, and 4. 6.) Which graph shows the quadratic function y = 3x2 + 12x + 14? The following graph is labeled A: A four quadrant graph with a parabola opening up, passing through the points (-3, 1), (-2, -2,) and (-1, 1) with the vertex at(2, -2). The following graph is labeled B: A four quadrant graph with a parabola opening up, passing through the points (1, 4), (2, 1,) and (3, 4) with the vertex at(2, 1). The following graph is labeled C: A four quadrant graph with a parabola opening up, passing through the points (-3, 5), (-2, 2,) and (-1, 5) with the vertex at(-2, 2). The following graph is labeled D: A four quadrant graph with a parabola opening up, passing through the points (1, 1), (2, -2,) and (3, 1) with the vertex at(2, -2). A) A B) B C) C D) D 9.) A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 34t + 80 The average rate of change of f(t) from t = 5 seconds to t = 7 seconds is _____ feet per second. 10.) What are the exact solutions of x2 + 5x + 1 = 0? A) x = the quantity of 5 plus or minus the square root of 21 all over 2 B) x = the quantity of negative 5 plus or minus the square root of 21 all over 2 C) x = the quantity of 5 plus or minus the square root of 29 all over 2 D) x = the quantity of negative 5 plus or minus the square root of 29 all over 2 12.) A company launched four new products. The market price, in dollars, of the four products after different number of years is shown below: Table displaying data. The first column one shows the heading number of years, and rows product 1, product 2, product 3, product 4. The next column heading says 1 with rows 2.00, 2.20, 2.20, and 2.80. The next column header says 2 with rows 4.00, 4.80, 6.40 and 3.20. The next column header says 3 with rows 8.00, 6.80, 11.60, 3.20. The last colunn heading says 4 with rows 16.00, 9.20, 18.80 and 2.80. Based on the data given, the price of which product will eventually exceed all others? A) Product 1 B) Product 2 C) Product 3 D) Product 4 13.)The following function represents the production cost f(x), in dollars, for x number of units produced by company 1: f(x) = 0.05x2 - 7x + 300 The following table represents the production cost g(x), in dollars, for x number of units produced by company 2: x g(x) 0.6 899.58 0.8 899.52 1 899.50 1.2 899.52 1.4 899.58 Based on the given information, the minimum production cost for company _____ is greater. [Put 1 or 2 in the blank space] 14.)A basketball is thrown upwards. The height f(t), in feet, of the basketball at time t, in seconds, is given by the following function: f(t) = -16t2 + 94t + 12 Which of the following is a reasonable domain of the graph of the function when the basketball falls from its maximum height to the ground? A) 3 < t < 6 B) 2 < t < 5 C) 1 < t < 4 D) 0 < t < 3 16.)A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? A) Add 9 to both sides of the equation B) Add 5 to both sides of the equation C) Subtract 5 from both sides of the equation D) Subtract 9 from both sides of the equation 17.)If (x + 1)2 = 55, what is the approximate positive value of x? A) 2.71 B) 5.42 C) 6.42 D) 7.35 19.) The following function shows the relationship between the selling price (s), and profit P(s), in dollars, for a company: P(s) = -20s2 + 1,400s - 12,000 Which statement best describes the intervals where the company's profit increases, decreases, or records a maximum? A) It is least when the selling price is $30. B) It is greatest when the selling price is $30. C) It decreases when the selling price increases beyond $35. D) It increases when the selling price increases beyond $35. 20.)The graphs of f(x) and g(x) are shown below: graph of function f of x open upward and has its vertex at negative 7, 0. Graph of function g of x opens upward and has its vertex at negative 7, negative 2. If f(x) = (x + 7)2, which of the following is g(x) based on the translation? A) g(x) = x2 + 18x + 81 B) g(x) = x2 + 10x + 25 C) g(x) = x2 + 14x + 47 D) g(x) = x2 + 14x + 51

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@AngelWilliams16

OpenStudy (dan815):

Hey these are too many questions.. one at a time, we here to help you learn.

OpenStudy (dan815):

Or you can work with someone that is also working on something like that, collaborate and learn

OpenStudy (anonymous):

Ok. Can you please help me with number 3?

OpenStudy (anonymous):

I attached a graph earlier

OpenStudy (anonymous):

3.) The graphs of f(x) and g(x) are shown below: graph of function f of x equals x squared minus x minus 12. Graph of function g of x equals negative 2.5 times x minus 2 What are the solutions to the equation f(x) = g(x)? A) x = -3, 4 B) x = -4, 2.5 C) x = -0.8, 2.5 D) x = 8, -8

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@SyedaLovesPie

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

@AngelWilliams16

OpenStudy (anonymous):

@victorhernandez

OpenStudy (anonymous):

LOOOOOL i would help you but im not that smart at stuff like that :3